Copper has one conduction electron per atom.Its density is 8.89 g/cc 4 its atomic mass is 63.54g/mol. If a Cu wire of diameter of 1nm carries a current of 2 A, what is the drift speed of electron in wire ?

Grade:Upto college level

1 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear Mainak ,

You Can use simple equation :

         I = neAvd   where n = no of conduction electrons /volume

                                       e = 1.6 x 10-19 c

                                       A = Area of Cross section of the wire

                                       vd = Drift velocity

calculation of n

  let us consider a seciton of length  1 m  , volume of that section  =  ∏ r2 l    = 7.85 *10-19  m3

         Mass of piece = ∏ r2 l * density  =  6.98  *10-12  g

    therefore no of moles in that amount  =  6.98  *10-12  g / 63.54 g  = 1.098 *10-13 

  no of atoms of copper in this amount  =  1.098 *10-13 * 6.022 *1023  = 6.617 * 1010  atoms  =  6.617 * 1010 electrons

there fore  , n  = 6.617 * 1010 electrons / 7.85 *10-19  m3 =  8.43  * 1028

 

now apply equations ,        2 = 8.43  * 1028 *1.6 *10-19 * 7.85 *10-19 *vd

                                            vd = 1.88 * 108  m/s 

the value comes out to be too high because the wire is of nano scale, it is still impossible to buildsuch wires.

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.